5^(x+7)=625^(x-7)
x= ?

Answer by orca(409)

Note that 625 = 5^4.
So the original equation 5^(x+7)=625^(x-7) becomes:
5^(x+7)=5^4(x-7)
Therefore x+7 = 4(x-7)
Solving the above equation for x, we have
x + 7 = 4x - 28
3x = 35
x = 35/3